# The Honest Truth About Causal Trees: Accuracy Limits for Heterogeneous Treatment Effect Estimation

### v1

@@ 6.1 Decision Stumps
The phenomenon of generating unbalanced cells in adaptive recursive partitioning schemes has been observed in various forms since the inception of CART. Historically, this phenomenon has been called the end-cut preference, where splits along noisy directions tend to concentrate along the boundary of the parent node. More specifically, considering the standard CART for regression estimation without sample splitting, Breiman et al. (1984, Theorem 11.1) and Ishwaran (2015, Theorem 4) showed that in one-dimension ($p=1$), for each $\delta\in(0,1)$, $\mathbb{P}(n(\mathsf{t}_{\mathtt{L}})\leq\delta n\;\text{or}\;n(\mathsf{t}_{\mathtt{R}})\geq(1-\delta)n)\to 1$ as $n\to\infty$. If applicable to the context of this paper, their result would only imply rates in uniform norm slower than any constant multiple of the already nearly optimal rate $\sqrt{n/\log\log(n)}$, i.e., for any $C>0$,
@@
In contrast, our results hold for all $p\geq 1$ and characterize precisely the regions of the support $\mathcal{X}$ where the pointwise rates of estimation are slower than any polynomial-in-$n$ (see Corollary SA-7, Theorem SA-14, Corollary SA-21 in the supplemental appendix). Thus, past theoretical work is not strong enough to illustrate the weaknesses of causal trees for pointwise estimation (i.e., prior lower bounds in the literature would be too loose to be informative). Furthermore, our results also study settings where sample splitting (honesty) is used, and demonstrate that they cannot mitigate the low convergence rate of adaptive causal trees under Assumption 1. Last but not least, our results apply to the causal tree constructions which are different (and more complicated) than those in plain vanilla CART regression (Definition 2).
@@ 6.2 Deeper Trees, Multivariate Covariates, and the Location of Small Cells
Our theoretical results show that, under Assumption 1, the first split of any decision tree construction will generate a small child cell with non-vanishing probability. As a result, and due to their recursive nature, deeper tree constructions will have multiple regions with too small sample sizes (with non-vanishing probability). This problem is exacerbated in multiple dimensions ($p>1$), which is exactly the setting where causal tree estimators would be potentially more useful to uncover treatment effect heterogeneity.
The small regions of the support $\mathcal{X}$, and hence the slower than any polynomial-in-$n$ convergence rate (or inconsistency) of causal tree estimators, need not occur near a region of the boundary of $\mathcal{X}$. At each stage in the tree construction, a parent node $\mathsf{t}$ will generate two child nodes, one small and the other large, but the splitting may realize anywhere on $\mathsf{t}$ (parent cell) and along any individual covariate (in $\mathbf{x}_{i}$, or axis), thereby generating problematic hyper-rectangle cells all over the support $\mathcal{X}$ with non-vanishing probability.
@@ 6.3 Regularization and Bias
It is tempting to try to regularize the decision tree estimator in order to eliminate the small cell problem, and thus improve its convergence rate. For instance, the tree construction algorithm may not split a parent node if the effective sample size is to small, or it may include a penalty term for overfitting. However, it is also important to note that adaptive decision tree constructions purposely select small cells for two opposing reasons: misspecification bias vs. low signal-to-noise ratio. More precisely, on the one hand, if the unknown conditional expectation function exhibits high curvature (bias) in a certain region of $\mathcal{X}$, then the tree construction will tend to generate a small child cell (node) in that region to reduce misspecification bias, which is precisely a celebrated feature of an “adaptive” procedure. On the other hand, as shown in this paper, small cells also emerge with non-vanishing probability when there is no misspecification bias in that region, that is, when the unknown conditional expectation function is locally constant. In practice, it is impossible to distinguish between the two equality possible scenarios.
Our theoretcal results purposely remove misspecification bias by considering data generating processes with constant conditional expectation functions. In real application settings, however, the conditional expectation functions may exhibit heterogeneity (even if locally constant), in which case regularization to remove small cells may led to large bias in the causal decision tree estimators, also affecting their convergence rate.
@@ 6.4 $\alpha$-Regularity and Causal Random Forests
Under specific assumptions, Wager and Athey (2018) and others established polynomial-in-$n$ convergence rates for honest causal trees and forests. The slow convergence rates establish in Theorem 1 do not contradict, but are rather precluded by existing polynomial-in-$n$ convergence guarantees in the literature because they assume that each split generates two child nodes that contain a constant fraction of the number of observations in the parent node, i.e., $n(\mathsf{t}_{\mathtt{L}})\gtrsim n(\mathsf{t})$ and $n(\mathsf{t}_{\mathtt{R}})\gtrsim n(\mathsf{t})$. The key assumption is often called $\alpha$-regularity, because it assumes that the tree construction generates an $\alpha>0$ proportion of the data in each terminal node (cell).
Our theoretical results imply that assumptions such as $\alpha$-regularity, or variants thereof, which require balanced cells almost surely, are incompatible with standard decision tree constructions employing causal trees (Athey and Imbens, 2016) or any other conventional CART methodology (e.g., Behr et al., 2022, and references therein). By implication, results for causal random forests relying on $\alpha$-regularity, or variants thereof, do not apply to standard recursive partitioning using CART-type algorithms. Some form of (algorithmic and/or statistical) regularization is needed, thereby introducing a bias in the estimation as well as additional tuning parameters that would need to chosen in practice.
@@ 6.5 Decision Tree Regression
The supplemental appendix also studies standard adaptive decision tree regression via CART for nonparametric estimation of the conditional expectation of an output given a collection of features. Section SA-2 in the supplemental appendix establishes an analogue of Theorem 1, demonstrating that adaptive decision tree regression exhibits slow convergence rate or inconsistency, as causal trees do, depending on the sample splitting design used.
[…]
6.6 Invalidity of Inference Methods
Theorem 1 establishes lower bounds on the uniform convergence rate of causal decision tree estimators. The main technical observation is that these estimation procedures will generate a partition of $\mathcal{X}$ with highly unbalanced cells, where potentially many cells will have a very small number of samples. These results are established under Assumption 1, which does not assume a parametric family of distributions on the data, but rather only independence and moment conditions.
From an inference perspective, our results also show that a valid (Gaussian or otherwise) distributional approximation for the causal decision tree estimators, after perhaps properly centering and scaling, does not hold in general. The main obstacle is that the effective sample size may not even increase for the approximation to apply in many regions of $\mathcal{X}$. In particular, standard inference methods, such as the usual confidence intervals of the form $\hat{\tau}_{l}^{q}(\mathbf{x})\pm z_{\alpha}\cdot\text{Sd.Err.}(\hat{\tau}_{l}^{q}(\mathbf{x}))$ with $z_{\alpha}$ denoting the usual quantile of the standad Gaussian distribution, $\text{Sd.Err.}()$ a standard error estimator, and $q\in\{\mathtt{NSS},\mathtt{HON},\mathtt{X}\}$, will not deliver asymptotically valid inference for the parameter of interest $\tau(\mathbf{x})$.
@@ 7 Simulations
We illustrate the implications of Theorem 1 in the univariate case $p=1$. Figure 1 reports the pointwise root mean squared error $\mathrm{RMSE}(x)\;=\;\bigl{\{}\mathbb{E}\bigl{[}(\hat{\tau}_{\ell}^{\,q}(x)-\tau)^{2}\bigr{]}\bigr{\}}^{1/2}$, for $\ell\in\{\mathtt{DIM},\mathtt{IPW},\mathtt{SSE}\}$ and $q\in\{\mathtt{NSS},\mathtt{HON},\mathtt{X}\}$, estimated from 2,000 Monte Carlo replications under $\tau=\mu_{0}=\mu_{1}=0$, $\varepsilon_{i}(0),\varepsilon_{i}(1)\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathsf{N}(0,1)$, $X_{i}\sim\mathrm{Uniform}[0,1]$, and n=1,000n=1{,}000. For each of the nine causal-tree estimators, we consider depths $K\in\{1,\dots,5\}$, where curves are color-coded by $K$.
Figure 1: Plots of root mean-squared error (RMSE) of heterogeneous treatment effect estimation using nine distinct causal tree methods with depth $K=1,2,\cdots,5$. We chose $p=1$, and the univariate covariate $X$ is supported on $[0,1]$. For all methods and depths, the causal tree has smallest pointwise RMSE near the center of the covariate space, but the performance degrades as the evaluation points move closer to the boundary. The experiment is conducted with 2,000 Monte-Carlo simulations.
@@
Two patterns emerge across all nine methods: (i) For any fixed $K$, the pointwise RMSE is smallest near the center of the covariate space and increases as $x$ approaches the boundary; (ii) For any fixed $x\in[0,1]$, the RMSE increases with tree depth $K$. The first pattern is due to the small cells near boundary predicted by (6), rendering a situation where local averaging is less accurate. The second is consistent with the $\mathtt{X}$-results of Theorem 1 and, heuristically, extends to $\mathtt{NSS}$ and $\mathtt{HON}$: at higher depths, a larger fraction of evaluation points lie near terminal node boundaries, where the same boundary effects that govern decision stumps degrade performance, leading to increased RMSE even for interior points.
@@ Acknowledgments
The authors thank Benjamin Budway, Max Farrell, Boris Hanin, Felix Hoefer, Michael Jansson, Joowon Klusowski, Boris Shigida, Jantje Sönksen, Jennifer Sun, Rocio Titiunik, and Kevin Zhang for comments. Cattaneo gratefully acknowledges financial support from the National Science Foundation through SES-1947805, SES-2019432, and SES-2241575. Klusowski gratefully acknowledges financial support from the National Science Foundation through CAREER DMS-2239448.

### v2

@@ 6 Discussion
This section interprets the theoretical implications of Theorems 13 and related results for adaptive regression trees established in the supplemental appendix.
@@ 6.1 Decision Stumps
The generation of highly unbalanced cells in adaptive recursive partitioning has been recognized since the early development of CART and is often referred to as the end-cut preference. Informally, when the signal is weak relative to sampling noise, the empirical splitting criterion may be optimized by thresholds located near the boundary of the parent node.
For example, in the context of standard CART regression without sample splitting, Breiman et al. (1984, Theorem 11.1) and Ishwaran (2015, Theorem 4) showed that in one-dimension ($p=1$), for each $\delta\in(0,1)$, $\mathbb{P}(n(\mathsf{t}_{\mathtt{L}})\leq\delta n\;\text{or}\;n(\mathsf{t}_{\mathtt{R}})\geq(1-\delta)n)\to 1$ as $n\to\infty$. If applicable in our setting, their result would only imply uniform convergence rates slower than any constant multiple of the already nearly optimal rate $\sqrt{n/\log\log(n)}$, i.e., for any $C>0$,
@@
In contrast, our results hold for all $p\geq 1$ and precisely characterize the regions of the support $\mathcal{X}$ where the pointwise rates of estimation are slower than any polynomial-in-$n$ (see Theorems SA-2 and SA-20, as well as Corollaries SA-10 and SA-30 in the supplemental appendix). Thus, existing theoretical results are not strong enough to reveal the limitations of causal trees for pointwise estimation. Moreover, our analysis covers settings with sample splitting (honesty) and shows that this modification does not mitigate the slow convergence of adaptive causal trees under Assumption 1. Finally, our results apply to causal tree constructions that differ from, and are more complex than, standard CART regression trees.
@@ 6.2 Deeper Trees, Multivariate Covariates, and the Location of Small Cells
Our theoretical results show that, under Assumption 1, the first split of an adaptive decision tree generates a small child cell with non-vanishing probability. Because recursive partitioning propagates this imbalance across subsequent levels of the tree, deeper trees necessarily produce multiple regions with very small effective sample sizes, again with non-vanishing probability. This phenomenon becomes more pronounced as the covariate dimension increases ($p>1$), precisely the regime in which tree-based methods are often employed to detect treatment effect heterogeneity.
Importantly, these problematic regions need not occur near the boundary of $\mathcal{X}$. At each stage of the recursive construction, a parent node $\mathsf{t}$ produces one large and one small child node, but the split can occur anywhere within $\mathsf{t}$ and along any coordinate direction. As a consequence, adaptive tree constructions can generate collections of hyper-rectangular cells with extremely small sample sizes scattered throughout the interior of the covariate space $\mathcal{X}\subseteq\mathbb{R}^{p}$.
@@ 6.3 Regularization and Bias
A natural response to the small-cell phenomenon is to regularize the tree construction in order to prevent highly imbalanced splits. For instance, the algorithm may impose minimum node size constraints or incorporate penalties designed to discourage overfitting. Such modifications can reduce estimation variance but necessarily alter the class of admissible partitions.
Adaptive tree procedures generate small cells for two distinct reasons. When the conditional expectation function exhibits strong local curvature, finer partitions reduce approximation bias and constitute a key advantage of adaptive methods. At the same time, as shown in this paper, highly imbalanced cells also arise with non-vanishing probability even when the regression function is locally flat. In practice, these two sources of small cells cannot be distinguished reliably.
Our theoretical analysis isolates the variance mechanism by focusing on data generating processes with constant conditional expectation functions. In empirical applications, however, treatment effects may vary across the covariate space, so regularization rules designed to eliminate small cells may introduce substantial approximation bias and thereby affect overall convergence rates.
Even when small or imbalanced cells appear undesirable from a purely statistical point of view, they may still play a useful role in recursive partitioning. In particular, Ishwaran (2015) argues that imbalanced splits (e.g., the aforementioned end-cut preference) can be beneficial because they preserve a relatively large sample in one descendant node after a poor earlier split. To illustrate, consider the threshold CATE function $\tau(\mathbf{x})=\mathds{1}(x_{1}\leq 1/2)$ and suppose the tree first makes a poor split on an irrelevant coordinate, say $x_{2}=0.05$. Although this split carries no information about the treatment effect, it is highly imbalanced. That is, one child contains only a small fraction of the sample, and the other contains the majority of the observations, including the entire signal region $\{x_{1}\leq 1/2\}$. The algorithm may therefore continue splitting within this large node and still recover the relevant threshold in $x_{1}$. In contrast, if one imposes a minimum node-size requirement, then an early poor split on an irrelevant coordinate must divide the sample much more evenly, thereby shrinking the sample size on every downstream branch and leaving less room for later correction. In this sense, some imbalanced splits are not merely an incidental by-product of adaptation, but rather a key part of the mechanism by which the algorithm corrects itself locally. Accordingly, a blanket regularization that suppresses small cells may remove not only splits that are needed to reduce misspecification bias, but also splits that help deep trees recover from earlier mistakes.
@@ 6.4 $\alpha$-Regularity and Causal Random Forests
Theorem 1 does not contradict existing (suboptimal) polynomial-in-$n$ convergence rate results for honest causal trees and forests (Wager and Athey, 2018). Rather, it shows that those results rely on strong algorithmic balance conditions that are incompatible with standard CART-type greedy splitting. In particular, prior analyses assume that each split allocates a fixed proportion of observations to both child nodes, that is, $n(\mathsf{t}_{\mathtt{L}})\geq\alpha\,n(\mathsf{t})$ and $n(\mathsf{t}_{\mathtt{R}})\geq\alpha\,n(\mathsf{t})$, where $\alpha\in(0,1/2]$. This assumption is often called $\alpha$-regularity, because it assumes that the tree construction generates an $\alpha$-proportion of the data in each child node (cell). Although $\alpha$-regularity is technically convenient and underlies positive consistency results for sufficiently deep trees and forest ensembles, it can force the resulting partition to be substantially less parsimonious than the underlying data generating mechanism would warrant, with corresponding losses in accuracy, adaptivity, and interpretability.
In particular, $\alpha$-regularity may substantially alter the adaptive behavior of recursive partitioning. To illustrate, consider again the threshold CATE function
$\tau(\mathbf{x})=\mathds{1}(x_{1}\leq\eta),\qquad\eta\in(0,1),$
with $\mathbf{x}=(x_{1},\ldots,x_{p})^{\top}$ uniformly distributed on $0,1]^{p}$. At the population level, an unconstrained CART procedure can recover this function with a single split on the relevant variable $x_{1}$ at $x_{1}=\eta$. In contrast, if $\eta<\alpha$, an $\alpha$-regular tree cannot make that split at the root. It must instead approximate the threshold through a sequence of progressively finer splits, requiring at least
$\left\lceil\log_{1/\alpha}(1/\eta)\right\rceil$
successive splits along $x_{1}$. In addition to reducing parsimony, it also weakens variable-selection adaptivity. Rather than isolating the relevant subgroup immediately by one decisive split on $x_{1}$, the procedure is forced to spend multiple splits refining the same coordinate. In finite samples, especially when $p$ is large relative to $n$, this repeated refinement creates more opportunities for spurious splits on irrelevant covariates.
More generally, $\alpha$-regularity limits the ability of the tree to isolate small but substantively meaningful regions of the covariate space. While sufficiently deep balanced trees may still approximate such structures in an integrated sense, the resulting partitions need not recover the underlying subgroup itself. In applications where subgroup identification is a primary objective, ensemble averaging (e.g., random forests) does not resolve this limitation.
It is also important to note that $\alpha$-regularity is primarily a theoretical device rather than a feature of standard applied implementations. In practice, regularization is typically governed by software defaults or light tuning, with minimum node-size choices fixed at small values (e.g. $2$, 55, 2020, or $30$) rather than scaled with the sample size. Consequently, convergence guarantees derived under balance conditions apply to a more constrained algorithm than the CART-type procedures commonly used in empirical work. Extending such guarantees to canonical implementations would require additional regularization that modifies the estimator and introduces further bias-variance trade-offs.
6.5 Invalidity of Inference Methods
Theorem 1 also has direct implications for statistical inference based on adaptive causal trees. Because recursive partitioning generates highly imbalanced cells with non-vanishing probability, the effective sample size within some regions of the covariate space need not increase with the overall sample size. As a consequence, standard distributional approximations for (“honest”) causal tree estimators may fail to hold even after appropriate centering and scaling. In particular, Gaussian asymptotic approximations can break down in regions where terminal nodes contain only a small number of observations.
This phenomenon undermines commonly used inference procedures based on asymptotic normality. For example, confidence intervals of the form
$\hat{\tau}_{l}^{q}(\mathbf{x})\pm z_{\alpha}\cdot\text{Sd.Err.}(\hat{\tau}_{l}^{q}(\mathbf{x})),$
where $z_{\alpha}$ denoting the usual quantile of the standard Gaussian distribution, $\text{Sd.Err.}()$ a standard error estimator, and $q\in\{\mathtt{NSS},\mathtt{HON},\mathtt{X}\}$, need not provide asymptotically valid coverage for $\tau(\mathbf{x})$ over many regions of $\mathcal{X}\subseteq\mathbb{R}^{p}$.
@@ 6.6 Decision Tree Regression
The supplemental appendix establishes analogous lower bounds for standard CART regression trees used for nonparametric estimation of conditional expectation functions. In particular, Section SA-2 shows that adaptive regression trees can exhibit similarly slow uniform convergence, and may even be inconsistent depending on the sample-splitting design.
These findings complement earlier large-sample analyses of the decision stump without sample splitting by Bühlmann and Yu (2002) and Banerjee and McKeague (2007). In the univariate setting ($p=1$, $K=1$), those studies showed that the empirical minimizers at the root node converge to well-defined population minimizers at cube-root rate under smoothness and identification conditions on the regression function.
Our analysis demonstrates that such conclusions are not uniformly valid over broad classes of regression functions. In particular, constant regression functions must be excluded from the admissible class for cube-root convergence results to hold uniformly over the support of the covariate. This distinction highlights the importance of uniform-in-function performance guarantees when evaluating adaptive recursive partitioning procedures.
@@ 7 Simulations
We illustrate the implications of Theorem 1 in the bivariate case ($p=2$). Figure 1 reports the pointwise root mean squared error $\mathrm{RMSE}(\mathbf{x})=\bigl\{\mathbb{E}\bigl[(\hat{\tau}_{\ell}^{\,q}(\mathbf{x})-\tau)^{2}\bigr]\bigr\}^{1/2}$, for $\ell\in\{\mathtt{DIM},\mathtt{IPW},\mathtt{SSE}\}$ and $q\in\{\mathtt{NSS},\mathtt{HON},\mathtt{X}\}$, estimated from $2,000$ Monte Carlo replications under $\tau=\mu_{0}=\mu_{1}=0$, $\varepsilon_{i}(0),\varepsilon_{i}(1)\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathsf{N}(0,1)$, $\mathbf{x}_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathrm{Uniform}([0,1]^{2})$, and $n=1,000$. For each of the nine causal tree estimators, we consider tree depths $K\in\{1,\dots,5\}$, where the curves are color-coded by $K$.
Two patterns emerge across all nine methods. First, for any fixed depth $K$, the pointwise RMSE is the smallest near the center of the covariate space and increases as $\mathbf{x}$ approaches the boundary. This pattern reflects the small-cell phenomenon predicted by (6): splits occurring near the boundary produce highly imbalanced nodes, reducing the effective sample size available for local averaging.
@@
Second, for any fixed $\mathbf{x}\in[0,1]^{2}$, the RMSE increases as the tree depth $K$ increases. This behavior is consistent with the $\mathtt{X}$-adaptive results in Theorem 1 and appears heuristically in the $\mathtt{NSS}$ and $\mathtt{HON}$ cases as well. As the depth increases, a larger fraction of evaluation points lie close to terminal node boundaries, where the same boundary effects that govern decision stumps lead to higher estimation variance. Consequently, the RMSE increases even for points located in the interior of the covariate space.
Figure 1: Root mean-squared error (RMSE) of heterogeneous treatment effect estimation using nine causal tree estimators with depth $K=1,2,\dots,5$. The covariate is bivariate ($p=2$) with support $0,1]^{2}$. Across all methods and depths, the smallest RMSE occurs near the center of the covariate space, while performance deteriorates as the evaluation point approaches the boundary. Results are based on $2,000$ Monte Carlo replications.
Appendix A Correction of Eicker (1979)
As part of the technical arguments used in this paper, we correct a statement concerning the limiting distribution of the maximum of an Ornstein-Uhlenbeck (O-U) process in Eicker (1979, Theorem 5). Specifically, the term $\log(c)$ should appear in place of $2\log(c)$, where $c>0$. This matters because replacing $\log(c)$ by $2\log(c)$ changes the factor $e^{-(z-\log(c))}=ce^{-z}$ to $e^{-(z-2\log(c))}=c^{2}e^{-z}$, so the limiting exponent depends quadratically rather than linearly on the window length. The problem is that this is incompatible with the Gaussian correlation inequality argument we used for adjacent windows. That is, for two adjacent windows of lengths $c_{1}\log n$ and $c_{2}\log n$, the event over the combined window is the intersection of the corresponding events on the two pieces, and the Gaussian correlation inequality implies that the probability for the combined window must be at least the product of the two individual window probabilities. With the correct term $\log(c)$, this is perfectly consistent, because the limiting probability for a window of length $(c_{1}+c_{2})\log n$ is
$\exp(-(c_{1}+c_{2})e^{-z})=\exp(-c_{1}e^{-z})\exp(-c_{2}e^{-z}),$
so the dependence on window length adds in exactly the way required by the inequality. If one instead used $2\log(c)$, then the same reasoning would force
$\exp(-(c_{1}+c_{2})^{2}e^{-z})\geq\exp(-(c_{1}^{2}+c_{2}^{2})e^{-z}),$
because the left-hand side would be the limiting probability for the combined window, while the right-hand side would come from the product lower bound for the two adjacent pieces. But this inequality is false, since $(c_{1}+c_{2})^{2}>c_{1}^{2}+c_{2}^{2}$ whenever $c_{1},c_{2}>0$. Thus the extra factor $2$ leads to a direct contradiction with the probability inequality implied by the Gaussian correlation argument. For completeness, we state below a corrected version of the result in a slightly more general form, allowing for the maximum of the norm of a possibly multivariate O–U process. Let $\Gamma(\cdot)$ denote Euler Gamma function.
Lemma 4 (Vector-Valued Markov-type Darling-Erdős).
Let $\{(V_{1}(t),\ldots,V_{d}(t)):0\leq t<\infty\}$ be $d$ independent identically distributed Ornstein-Uhlenbeck processes with $\mathbb{E}[V_{i}(t)]=0$ and $\mathbb{E}[V_{i}(t)V_{i}(s)]=\exp(-|t-s|/2)$, $1\leq i\leq d$. Define
$N(t)=\bigg(\sum_{1\leq i\leq d}V_{i}^{2}(t)\bigg)^{1/2}.$
For any $c>0$, $z\in\mathbb{R}$,
$\lim_{n\rightarrow\infty}\mathbb{P}\Big(a(\log(n))\sup_{0\leq t\leq c\log(n)}N(t)-b_{d}(\log(n))\leq z\Big)=\exp\Big(-e^{-(z-\log(c))}\Big),$
where $a(t)=(2\log(t))^{1/2}$ and $b_{d}(t)=2\log(t)+\frac{d}{2}\log\log(t)-\log\Gamma(d/2)$.
The proof of this result is given in the supplemental appendix (Section 4.1).
@@ Acknowledgments
The authors thank Benjamin Budway, Max Farrell, Boris Hanin, Felix Hoefer, Michael Jansson, Joowon Klusowski, Boris Shigida, Jantje Sönksen, Jennifer Sun, Rocio Titiunik, and Kevin Zhang for comments.
Funding
Cattaneo gratefully acknowledges support from the National Science Foundation through grants SES-2019432, DMS-2210561, SES-2241575, and SES-2342226 and from the National Institute for Food and Agriculture (NIFA) through grant 2024-67023-42704. Klusowski gratefully acknowledges support from the National Science Foundation through NSF CAREER grant DMS-2239448 and from the Alfred P. Sloan Foundation through a Sloan Research Fellowship.